OpenStudy (buttermequeasy):

An architect needs to determine the slope between two points on a ski lift. The two points have been identified as (15, 45) and (150, 60), where x is the horizontal distance and y is the vertical distance from the bottom of the lift. Assuming the lift runs in a straight line, what is the slope of the line between the two points? Help please! Urgent!

hero (hero):

Use the slope formula

hero (hero):

(60 - 45)/(150 - 15) = 15/135 = 1/9